The no. of points having both co ordinates as integers that lie in interior of triangle with vetices (0,0),(0,41),(41,0) ???
 

Grade:12

1 Answers

Vikas TU
6 years ago
 This is a privilege calculated triangle with right point at source 
presently consider (0,0) ,(0,1),(0,2) ...(0,41) we have 41 focuses 
next consider(1,0),(1,1),(1,2) ...(1,40) we have 40 focuses 
so no of focuses inside the triangle and on limit together are 
41+40+39+38+......1 
so add up to no of focuses =(41)(42)/2 = 861 
presently we have 81 focuses on limit 
therfore reqd no of focuses = 861-80= 780